if the length and width of a rectangle are doubled, by what factor is the length of its diagonal multiplied?
Answer by stanbon(75887)

if the length and width of a rectangle are doubled, by what factor is the length of its diagonal multiplied?
-----
If length = "L" and width = "W", diagonal = sqrt[L^2+W^2]
----
Doubling:
If length = "2L" and width = "2W",
diagonal = sqrt((2L)^2 + (2W)^2) = sqrt(4(L^2+W^2)) = 2*sqrt(L^2+W^2)
-----
Ans: by a factor of "2".
Cheers,
Stan H.
==============